OpenStudy (anonymous):

how do i solve this problem: (g * f)(x) = g(f(x)) the multiplication symbol(*) is an open circle g(x)=3x+1 f(x)= x/2

OpenStudy (amistre64):

(g o f)(x) is simple enough to read and then you dont need to "define" what the operation is spose to be

OpenStudy (amistre64):

place f(x) into g(x) at every place that you see an x

OpenStudy (anonymous):

is the answer 3(x/2)+1

OpenStudy (amistre64):

g(x)=3x+1 g(f(x)) = 3(f(x)) + 1 and restate f(x) with its rule; f(x) = x/2, therefore g(f(x)) = g(x/2) = 3(x/2) + 1

OpenStudy (amistre64):

you might need to "simplify" it further, but im never that good with knowing how simplified a simplified form is spose to go